# Explainable AI needs formalization

### v4

@@ Introduction
The use of machine learning (ML) holds great promise in many fields, including high-risk domains such as medicine. Regulations like the European AI Act demand that “high-risk AI systems shall be designed and developed […] to enable deployers to interpret the system’s output and use it appropriately” [28]. This need for “human-understandable” descriptions of the functions implemented by individual ML models is seemingly addressed by the field of “explainable artificial intelligence” (XAI). However, the formal basis of XAI is underdeveloped. Consequently, the possibility of using XAI for ML quality assurance is currently strongly limited.
@@
An example would be a neural network predicting the clinical outcome for patients in critical care from clinical and demographic patient characteristics. Here, different characteristics such as age or the presence of pre-existing conditions correspond to individual input features $x_{i}$, while the outcome of interest, such as death, corresponds to the target $y$. The neural network $f_{\bm{\theta}}$ learns the mapping between inputs and targets, where $\bm{\theta}$} represents the learnable parameters of the network.
@@
“Explainable Artificial Intelligence” (XAI) is an umbrella term for algorithms aiming to provide insight into the properties of ML models, their training data, a given test input submitted to the model, and/or the interplay between these. The predominant XAI paradigm is feature attribution, which refers to attributing an “importance” score $e_{i}$ to each input feature $x_{i}$. A distinction is made between global methods, where the attribution ${\mathbf{e}}=[e_{1},\ldots,e_{q}]^{\top}$ is a property of the model only, and local methods, where the attribution ${\mathbf{e}}({\mathbf{x}})=[e_{1}({\mathbf{x}}),\ldots,e_{q}({\mathbf{x}})]^{\top}$ additionally depends on the input. For the example in described in Supervised Machine Learning, global methods would assign each predictor, such as age, a constant importance score, whereas local methods would assign a score specific to each patient.
@@
One may also be interested in whether a model “bases” its decisions on confounding variables. Confounders induce correlations between training in- and outputs that can be used by the model for prediction. This can be problematic if the same correlations are not present in a testing context, leading the model to perform poorly. Lapuschkin et al. [51] study a case in which a watermark in images indicates such confounding, and use XAI methods in the process of identifying this effect.
@@
It is frequently assumed that XAI could be used to identify features, the manipulation of which would change a model’s output, a task also known as algorithmic recourse. For example [see 78], a bank might use an ML model to predict the return probability of a loan. For a known model and a given input, XAI would then be able to recommend changes of input variables (e.g., ‘increase salary’) to turn a negative outcome into a positive one. Similarly, it is assumed that XAI can help to verify that protected attributes (e.g., gender, race) do not influence model decisions. In an intensive care unit, an ML model might be used to predict mortality or other severe outcomes. Using XAI to identify possible intervention targets, such as medications, in this context [e.g., 6] implies that interventions have real-world consequences on the target variable beyond just changing the model output.
@@
In the following, we argue that the use of feature attributions for the above-mentioned explanation purposes amounts to asserting that the SAP holds; in other words, the SAP is a necessary property for XAI methods to serve these purposes. Subsequently, we discuss results presented in Haufe et al. [35, 50, 87], showing that a wide range of popular local and global feature attribution methods in fact do not possess the SAP, thus prohibiting conclusions about associational or even causal relations between features and target on the basis of these methods. As a consequence, these methods also fail to reliably serve the purposes mentioned in Desired purposes of XAI.
@@
Features like $X_{2}$ in Examples A and B, which improve predictions without being predictive themselves, are called suppressor variables in causal terminology [23]. Causal diagrams [see 61] of the generative models in both examples are provided in Figure 1 (c). Broadly speaking, any variable that is not informative (statistically associated with the target) itself but statistically related to an informative variable (e.g., modulating it through an independent mechanism) is a suppressor. Suppressors occur widely in real-world datasets and hamper model interpretation. As one example, the prevalence of a disease may be related to a person’s blood pressure but not their age. However, as blood pressure has an age-dependent baseline, the model might need to adjust its prediction with respect to that baseline in order to remove irrelevant variance introduced by age. Age, thereby, becomes a suppressor variable. In image classification, non-discriminative features such as lighting or weather conditions, or non-discriminative objects occluding class-specific objects, could be suppressors.
@@
Since suppressor variables have no statistical or causal association with the target variable, suppressor attributors do not possess the SAP, which has implications regarding their expected utility for the purposes introduced in Desired purposes of XAI. For example, suppressor features may often not coincide with prior expectations of an expert. Therefore, suppressor attributors cannot be used in a straightforward way to validate the correctness and fitness-for-purpose of models or their decisions using expert knowledge as insinuated by Ribeiro et al. [63]. Moreover, since it cannot be concluded that the highlighted features are part of previously unknown interactions or are causally related to the output, these methods cannot be reliably used to facilitate scientific discoveries or to identify incorrect models. For example, high importance on a protected attribute does not necessarily mean that the method “uses” this attribute for prediction. The model may also just remove variance related to that attribute from other informative variables. Finally, a prerequisite for identifying confounding variables causally influencing both in- and outputs of a model is to be able to recognize features with a statistical association to the target in the first place. The inability of suppressor attributors to distinguish such features from suppressor variables, as discussed here, thus implies that XAI methods cannot answer causal questions, such as questions related to confounding.
In both examples, any intervention on $X_{2}$ (through $X_{2}$ or $H$) would have no effect on $Y$ in the real world. In Example B, it would not even affect the model output, as the model is invariant to changes in $X_{2}$ by construction. In Example A, possible interventions on $X_{2}$ through $H$ could affect the model output; however, not in ways that would correlate with changes in $Y$.
@@
Faithfulness [Pixel Flipping, 66]
@@
SHAP [Marginal Expectation, 52]
@@
The results presented above have been established through joint theoretical analyses of data-generating processes, ML models, and feature attribution methods as well as through simulations using synthetic data with known ground-truth explanations. These techniques are not currently part of the standard toolkit for assessing the quality of explanations and XAI methods, pointing to the following fundamental structural limitations in the way the field assesses itself.
@@
Existing theoretical work has postulated axioms that are desirable for XAI methods to fulfill. For example, according to Sundararajan et al. [75], a method satisfies sensitivity, if a) for every input and baseline that differ in one feature but have different predictions, the differing feature is given non-zero importance, and if also b) the importance of a variable is always zero if the function implemented by the deep network does not depend (mathematically) on it. Axioms like this encode meaningful sanity checks but do not provide a notion of correctness or utility-for-purpose of an explanation. Several authors have proposed to close this gap by describing criteria for the “faithfulness” or “fidelity” of XAI methods. These concepts, however, are often not formulated in mathematically stringent form [see, 33, 43]. Moreover, faithfulness is insufficient to serve the purposes mentioned in Desired purposes of XAI, as we note further below.
@@
Various authors [e.g., 20, 65] make a distinction between “explainable AI”, which would include post-hoc feature attribution methods, and “interpretable AI”, which would include model architectures that are presumed to be intrinsically understandable to humans due to their simplicity. The latter are also occasionally referred to as “glassbox” models [62], and examples include linear models, generalized additive models (GAMs), models with sparse coefficients, and decision trees. However, what concrete interpretations such models are thought to afford is rarely stated. In the above Examples A and B, the Bayes-optimal linear models are uniquely defined and assign non-zero weights to suppressor variables, prohibiting certain desired interpretations and precluding certain actionable consequences, as demonstrated in Haufe et al. [35] and Wilming et al. [87]. Analogous fallacies apply to GAMs as shown by Clark et al. [22]. These works highlight that trained ML models, no matter how simple their structure, cannot be univocally interpreted without knowledge of the causal structure of their training data. The standard interpretation of models such as linear models and GAMs implicitly assumes statistically independent features, thereby sharing a fundamental limitation with post-hoc feature attribution methods.
@@
Existing frameworks for empirical XAI evaluation [e.g., 37] often primarily focus on secondary desiderata such as robustness of explanations instead of providing quantifiable notions of correctness. Nevertheless, “faithfulness” metrics are widely considered to be suitable surrogates for assessing explanation correctness. The most widely adopted operationalization of faithfulness is that the ablation (e.g., omission or obfuscation) of an important feature will lead to a drop in a model’s prediction performance. The presence of such a drop is then used to assess “correctness”. Popular perturbation approaches include permutation feature importance [17], stability selection [53], pixel flipping [66], RemOve And Retrain [ROAR, 39], and Remove and Debias [ROAD, 64], and prediction difference analysis [e.g., 13]. A variation is the model parameter randomization test [MPRT, 2].
@@
Several studies [e.g., 27, 38, 11] consider human judgment for XAI validation, where human experts either annotate inputs ex ante to provide ground-truth explanations or are asked to judge the quality of explanations ex post. While important, such approaches are insufficient as (sole) validations due to the possibility of both Type-I and Type-II errors in human judgments. For example, there may be complex statistical patterns in the data that are leveraged by ML models but are (currently) unknown to humans. This may lead an expert to reject a correct explanation. Human-computer interaction studies are considered an objective way to quantify the added value of AI explanations by some authors [e.g., 45]. Such studies compare the joint performance of a human user with access to an XAI with the performance of the user knowing only the outcome of the AI’s prediction, the performance of the user alone, and the performance of the AI alone. However, there are a growing number of studies reporting no correlation between the presence of explanations and combined human-XAI task performance, no correlation between explanation-based human prediction of AI performance and actual AI performance, and no correlation between explanation-induced human trust in AI decisions and actual AI performance [18, 10]. These results speak to the presence of a variety of human biases and psychological factors that hamper attempts to objectively evaluate XAI methods using human judgment alone. In fact, Bansal et al. [10] find that ‘humans will accept the AI’s recommendation, regardless of its correctness’, while Trout [77] discusses how human cognitive biases can generally lead to a wrong sense of understanding incorrect explanations. Notably, overconfidence in XAI explanations can lead to circular reinforcement of wrong beliefs, whereby humans may adapt their judgment to incorrect explanations, ultimately harming scientific knowledge discovery and theory building. As an example, consider a model using an uninformative suppressor feature to remove non-discriminative variance from a target-informative feature. Since this suppression relationship is stable, an XAI method may consistently highlight the suppressor as being important for the prediction. Without further information about the role of the suppressor in the model, this may lead the receiver of an explanation to erroneously conclude that the suppressor carries indispensible discriminative information.
@@
XAI methods have been criticized in many further ways [e.g., 31, 74, 83, 30]. For example, the low robustness and consistency of XAI explanations has been noted [7]. Moreover, explanations provided by different XAI methods are often found to be inconsistent. This can be used by an adversary (e.g., the provider of an ML algorithm in need to explain a decision to a user) to provide arbitrary explanations [15]. Similarly, a wealth of quality metrics is available to measure properties such as faithfulness which are observed to be inconsistent in their ranking of XAI methods [36]. It has been noted that developers of XAI methods could present their own method as being particularly faithful by optimizing the choice of metric [14]. It has also been pointed out that XAI methods can be manipulated to yield arbitrary explanations [26, 89]. In image prediction tasks, XAI explanations are frequently observed to resemble results of simple edge detection filters [e.g., 2, 49, 21]. Many XAI methods also come in multiple variants, and the criteria for choosing methods and their hyperparameters are often not well justified or documented.
The fundamental limitation of the field, though, is the lack of formal specifications of XAI problems. To ensure the fitness of XAI methods for their intended purposes by design, we argue that the current paradigm of algorithm-first development should be replaced by a requirement-driven XAI development and validation process [see also 25]. We propose that such a process should consist of six steps:
@@
It is unreasonable to call a mapping from input features to real numbers an explanation without endowing these numbers with a well-defined formal interpretation [e.g. 56]. As indicated above, different stakeholders, such as ML developers, users (e.g., physicians or patients), and regulators, may intend to use XAI for different purposes associated with different information needs. These needs may concern properties of a given ML model, its training data, a given test input, or combinations of these, and may differ between use cases.
@@
To enable the targeted development of XAI methods tailored to specific purposes, informal information needs communicated by stakeholders need to be translated into formal specifications and requirements, which will inevitably lead to distinct XAI problems. In that sense, “explanation” is understood here as an umbrella term describing the provision of information to stakeholders. We acknowledge that there can be multiple distinct notions of explanation, and thus explanation correctness, depending on the information requested by stakeholders and provided by XAI. We note that, as the correctness of a formalization cannot itself be formally validated, it is critical to ensure that it matches stakeholder intentions. To this end, the assessment of stakeholder needs and their subsequent formalization should go hand in hand and is best carried out in inter-disciplinary and inter-professional teams.
An example for successful XAI problem formalization is the work of Karimi et al. [47] on minimal algorithmic recourse, defined as the identification of minimal interventions in model inputs that lead to desired changes in model output. Karimi et al. [47] formulate this problem as a constrained optimization problem over interventions in a given structural causal model (SCM). Through this formalization, they show that conventional counterfactual explanations [80, 78] provide solutions that may violate the causal structure of the data, thus representing infeasible or ineffective interventions. For example, such explanations may erroneously suggest interventions on suppressor variables as also pointed out by Wilming et al. [87]. More generally, previous work[85, 86, 16] proposes the SAP as a necessary, though not sufficient, requirement for important features in the context of the explanation purposes discussed here.
@@
Given formal problem or requirement specifications, it becomes possible to theoretically analyze existing XAI algorithms with respect to established formal criteria. This has led various authors to identify systematic failure modes of existing XAI methods [e.g., 71, 12]. Kindermans et al. [50] and Wilming et al. [87] analyzed popular feature attribution methods and found that many do not fulfill the SAP property in general. Such analyses can help to identify theoretical shortcomings and guide the development of novel, improved methods.
Similarly, it may be possible to devise algorithms that meet formal criteria. Karimi et al. [48, 47] present algorithms for solving the optimization problem underlying minimal algorithmic recourse for different classes of data-generating processes and different degrees of prior knowledge. Likewise, the Pattern approach [35] relates a fitted linear model univariately to each individual input feature, thereby avoiding possible misinterpretations due to correlated features. Pattern can consequently be shown to correctly reject suppressor variables in the studied Examples A and B. Recent generalizations such as PatternGAM [22] and PatternLocal [32] extend the Pattern concept to non-linear models and have shown performance gains in empirical benchmarks involving non-linear data [21]. These feature attribution methods fulfill the SAP under well-defined conditions, thereby ensuring that features with high attribution represent sensible starting points, if not solutions, for a variety of popular explanation goals.
@@
Once methods that are theoretically and/or empirically validated with respect to given XAI problems or goals are available, it becomes of interest to theoretically analyze, quantitatively benchmark, and algorithmically improve these with respect to secondary quality indicators such as robustness, fairness, and uncertainty calibration. To this end, dedicated benchmarking frameworks such as Quantus [37] can be of use. Moreover, it is crucial to present explanations in ways that are aligned with human cognition and social norms [54]. Finally, it can be worthwhile to expand the range of applicability of validated XAI methods. Along these lines, recent work has extended the concept of activation patterns [35] to non-linear and local explanation domains [22, 32].
@@
Just as ML in general, the field of XAI is fast-paced with clever novel methodological developments and empirical validation approaches being introduced each year. Recent advancements in algorithmic recourse [48], confounder detection [44], and generative modeling [40, 72] promise to address some of the limitations presented here. The systematic formalization and scrutinization of the field of XAI is a wider effort that will eventually make it possible to objectively assess the ability of approaches to solve specific XAI problems. This may lead to XAI-based workflows that can indeed be used to systematically perform quality assurance for ML – and that may eventually find their way into ML production processes and industry standards.

### v5

@@ Introduction
The use of machine learning (ML) holds great promise in many fields, including high-risk domains such as medicine. Regulations like the European AI Act demand that “high-risk AI systems shall be designed and developed […] to enable deployers to interpret the system’s output and use it appropriately” [32]. This need for “human-understandable” descriptions of ML models is seemingly addressed by the field of “explainable artificial intelligence” (XAI). However, this paper show that the current algorithm-first paradigm of XAI development has led to methods whose results are commonly misinterpreted, and whose ability to serve purposes commonly associated with XAI, such as quality assurance for ML, is thereby compromised. To overcome biased or confounded XAI results, the fields requires formalization that goes beyond current axioms and desiderata, which mostly spare out the most important issue of explanation correctness. XAI methods should be developed to solve well-defined problems following a process that i) defines appropriate formal criteria of XAI correctness and fitness for relevant purposes and, ii) employs theory and empirical validation based on ground-truth data to assess a given method’s adherence to these criteria. This paper further argues that XAI methods targeted at explaining the model function only are insufficient to address desired downstream purposes such as diagnosing and correcting data and models, scientific discovery, and identifying intervention targets.
@@
An example would be a neural network predicting the clinical outcome for patients in critical care from clinical and demographic patient characteristics. Here, different characteristics such as age or the presence of pre-existing conditions correspond to individual input features $x_{i}$, while the outcome of interest, such as death, corresponds to the target $y$. The neural network $f_{\bm{\theta}}$ learns the mapping between inputs and targets, where $\bm{\theta}$} represents the learnable parameters of the network. An introduction into supervised ML can be found in [15].
@@
“Explainable Artificial Intelligence” (XAI) is an umbrella term for algorithms aiming to provide insight into the properties of ML models, their training data, a given test input submitted to the model, and/or the interplay between these. The predominant XAI paradigm is feature attribution, which refers to attributing an “importance” score $e_{i}$ to each input feature $x_{i}$. A distinction is made between global methods, where the attribution ${\mathbf{e}}=[e_{1},\ldots,e_{q}]^{\top}$ is a property of the model only, and local methods, where the attribution ${\mathbf{e}}({\mathbf{x}})=[e_{1}({\mathbf{x}}),\ldots,e_{q}({\mathbf{x}})]^{\top}$ additionally depends on the input. For the example in described in Supervised Machine Learning, global methods would assign each predictor, such as age, a constant importance score, whereas local methods would assign a score specific to each patient. Recent reviews of predominant XAI paradigms, approaches, and associated claims can be found in [83, 60]. A comprehensive social science perspective on XAI is added in [59].
@@
In a similar vein, it is expected that XAI can help to verify that protected attributes (e.g., gender, race) do not influence model decisions.
One may also be interested in whether a model bases its decisions on confounding variables. Confounders induce correlations between training in- and outputs that can be used by the model for prediction. This can be problematic if the same correlations are not present in a testing context, leading the model to perform poorly. Lapuschkin et al. [55] study a case in which a watermark in images indicates such confounding, and use XAI methods in the process of identifying this effect.
@@
XAI is frequently used to identify features, the manipulation of which would change a model’s output, a task also known as algorithmic recourse. For example [see 85], a bank might use an ML model to predict the return probability of a loan. For a known model and a given input, XAI would then be able to recommend changes of input variables (e.g., ‘increase salary’) to turn a negative outcome into a positive one. It is typically assumed that such interventions would not only affect the model output but also the target variable the model was trained to predict (e.g., the probability of default). In an intensive care unit, an ML model might be used to predict mortality or other severe outcomes. Using XAI to identify possible intervention targets, such as medications, in this context [e.g., 8] implies that interventions have real-world consequences on the target variable beyond just changing the model output.
@@
In the following, we discuss results presented in Haufe et al. [39, 54, 94], showing that a wide range of popular local and global feature attribution methods in fact do not possess the SAP, thus prohibiting conclusions about associational or even causal relations between features and target on the basis of these methods.
In Implications of suppressor attribution on explanation purposes, we further argue that the use of feature attributions for all of the above-mentioned explanation purposes amounts to asserting that the SAP holds; in other words, the SAP is a necessary property for XAI methods to serve these purposes. As a consequence, the discussed methods fall short of reliably serving the purposes mentioned in Desired purposes of XAI.
@@
Features like $X_{2}$ in Examples A and B, which improve predictions without being predictive themselves, are called suppressor variables in causal terminology [27]. Causal diagrams [see 67] of the generative models in both examples are provided in Figure 1 (c). Broadly speaking, any variable that is not informative (statistically associated with the target) itself but statistically related to an informative variable (e.g., modulating it through an independent mechanism) is a suppressor. Suppressors occur widely in real-world datasets and hamper model interpretation. As one example, the prevalence of a disease may be related to a person’s blood pressure but not their age. However, as blood pressure has an age-dependent baseline, the model might need to adjust its prediction with respect to that baseline in order to remove irrelevant variance introduced by age. Age, thereby, becomes a suppressor variable. In image classification, non-discriminative features such as lighting or weather conditions, or non-discriminative objects occluding class-specific objects, could be suppressors.
@@
Implications of suppressor attribution on explanation purposes
Since suppressor variables have no statistical or causal association with the target variable, suppressor attributors do not possess the SAP, which has implications regarding their expected utility for the purposes introduced in Desired purposes of XAI. For example, suppressor features may often not coincide with prior expectations of an expert. Therefore, suppressor attributors cannot be used in a straightforward way to validate or invalidate the correctness and fitness-for-purpose of models or their decisions using expert knowledge as anticipated by Ribeiro et al. [69]. Similarly, high importance on a protected attribute does not necessarily mean that the method “uses” this attribute for prediction. The model may also just remove variance related to that attribute from other informative variables. Moreover, since it cannot be concluded that the highlighted features are part of previously unknown interactions or are causally related to the output, these methods cannot be reliably used to facilitate scientific discoveries. While counterfactual and algorithmic recourse techniques are able to point to changes in inputs that have desirable effects on model outputs, such counterfactuals should not be expected to have an equivalent effect on the real-world target themselves. As optimal models may rely on causal effects of the target as well as on suppressor variables, counterfactual explanations may also typically include such variables, the manipulation of which can have no effect on the target in the real world. Recommendations based on such counterfactuals, thus, bear relevance only for the purpose of changing the model output and would be irresponsible to apply to the target itself. For example, a medical condition manifesting in a certain symptom (e.g. a genetic condition affecting behavior) does not imply that the treatment of the symptom will also affect the condition even though the value of the treatment variable affects the model output. Finally, a prerequisite for identifying confounding variables causally influencing both in- and outputs of a model is to be able to recognize features with a statistical association to the target in the first place. The inability of suppressor attributors to distinguish such features from suppressor variables, as discussed here, thus implies that XAI methods cannot answer causal questions, such as questions related to confounding.
In both examples, any intervention on $X_{1}$, $X_{2}$, or $H$ would have no effect on $Y$ in the real world. In Example B, interventions on $X_{2}$ would not even affect the model output, as the model is invariant to changes in $X_{2}$ by construction. In Example A, possible interventions on $X_{2}$ through $H$ could affect the model output; however, not in ways that would correlate with changes in $Y$.
@@
Faithfulness [Pixel Flipping, 72]
@@
SHAP [Marginal Expectation, 56]
@@
The SAP presents an example of a formal criterion (here, a necessary condition) for XAI fitness for a variety of downstream purposes. Through joint theoretical analyses of data-generating processes, ML models, and feature attribution methods as well as through simulations using synthetic data with known ground-truth explanations, it was further demonstrated that the SAP presents a testable property that enables judgments of the correctness of specific methods for these purposes. These techniques are, however, not currently part of the standard toolkit for assessing the quality of explanations and XAI methods, pointing to the following fundamental structural limitations in the way the field assesses itself.
@@
Existing theoretical work has postulated axioms that are desirable for XAI methods to fulfill. For example, according to Sundararajan et al. [81], a method satisfies sensitivity, if a) for every input and baseline that differ in one feature but have different predictions, the differing feature is given non-zero importance, and if also b) the importance of a variable is always zero if the function implemented by the deep network does not depend (mathematically) on it. Axioms like this encode meaningful sanity checks but do not provide a notion of correctness or utility-for-purpose of an explanation. Several authors have proposed to close this gap by describing criteria for the “faithfulness” or “fidelity” of XAI methods. These concepts, however, are often not formulated in mathematically stringent form [see, 37, 47]. Moreover, faithfulness is insufficient to serve the purposes mentioned in Desired purposes of XAI, as we note further below.
@@
Various authors [e.g., 23, 71] make a distinction between “explainable AI”, which would include post-hoc feature attribution methods, and “interpretable AI”, which would include model architectures that are presumed to be intrinsically understandable to humans due to their simplicity. The latter are also occasionally referred to as “glassbox” models [68], and examples include linear models, generalized additive models (GAMs), models with sparse coefficients, and decision trees. However, what concrete interpretations such models are thought to afford is rarely stated. In the above Examples A and B, the Bayes-optimal linear models are uniquely defined and assign non-zero weights to suppressor variables, prohibiting certain desired interpretations and precluding certain actionable consequences, as demonstrated in Haufe et al. [39] and Wilming et al. [94]. Analogous fallacies apply to GAMs as shown by Clark et al. [26]. These works highlight that trained ML models, no matter how simple their structure, cannot be univocally interpreted without knowledge of the causal structure of their training data. The standard interpretation of models such as linear models and GAMs implicitly assumes statistically independent features, thereby sharing a fundamental limitation with post-hoc feature attribution methods.
@@
Existing frameworks for empirical XAI evaluation [e.g., 41] often primarily focus on secondary desiderata such as robustness of explanations instead of providing quantifiable notions of correctness. Nevertheless, “faithfulness” metrics are widely considered to be suitable surrogates for assessing explanation correctness. The most widely adopted operationalization of faithfulness is that the ablation (e.g., omission or obfuscation) of an important feature will lead to a drop in a model’s prediction performance. The presence of such a drop is then used to assess “correctness”. Popular perturbation approaches include permutation feature importance [20], stability selection [58], pixel flipping [72], RemOve And Retrain [ROAR, 43], and Remove and Debias [ROAD, 70], and prediction difference analysis [e.g., 16]. A variation is the model parameter randomization test [MPRT, 2].
@@
Several studies [e.g., 31, 42, 13] consider human judgment for XAI validation, where human experts either annotate inputs ex ante to provide ground-truth explanations or are asked to judge the quality of explanations ex post. While important, such approaches are insufficient as (sole) validations due to the possibility of both Type-I and Type-II errors in human judgments. For example, there may be complex statistical patterns in the data that are leveraged by ML models but are (currently) unknown to humans. This may lead an expert to reject a correct explanation. Human-computer interaction studies are considered an objective way to quantify the added value of AI explanations by some authors [e.g., 49]. Such studies compare the joint performance of a human user with access to an XAI with the performance of the user knowing only the outcome of the AI’s prediction, the performance of the user alone, and the performance of the AI alone. However, there are a growing number of studies reporting no correlation between the presence of explanations and combined human-XAI task performance, no correlation between explanation-based human prediction of AI performance and actual AI performance, and no correlation between explanation-induced human trust in AI decisions and actual AI performance [21, 12]. These results speak to the presence of a variety of human biases and psychological factors that hamper attempts to objectively evaluate XAI methods using human judgment alone. In fact, Bansal et al. [12] find that ‘humans will accept the AI’s recommendation, regardless of its correctness’, while Trout [84] discusses how human cognitive biases can generally lead to a wrong sense of understanding incorrect explanations. Notably, overconfidence in XAI explanations can lead to circular reinforcement of wrong beliefs, whereby humans may adapt their judgment to incorrect explanations, ultimately harming scientific knowledge discovery and theory building. As an example, consider a model using an uninformative suppressor feature to remove non-discriminative variance from a target-informative feature. Since this suppression relationship is stable, an XAI method may consistently highlight the suppressor as being important for the prediction. Without further information about the role of the suppressor in the model, this may lead the receiver of an explanation to erroneously conclude that the suppressor carries indispensible discriminative information.
@@
Besides suppressor attribution, XAI methods have been criticized in many further ways [e.g., 35, 80, 90, 34, 3]. For example, the low robustness and consistency of XAI explanations has been noted [9]. Moreover, explanations provided by different XAI methods are often found to be inconsistent. This can be used by an adversary (e.g., the provider of an ML algorithm in need to explain a decision to a user) to provide arbitrary explanations [18]. Similarly, a wealth of quality metrics is available to measure properties such as faithfulness which are observed to be inconsistent in their ranking of XAI methods [40]. It has been noted that developers of XAI methods could present their own method as being particularly faithful by optimizing the choice of metric [17]. It has also been pointed out that XAI methods can be manipulated to yield arbitrary explanations [30, 96]. In image prediction tasks, XAI explanations are frequently observed to resemble results of simple edge detection filters [e.g., 2, 53, 25, 24]. As the results of a controlled study, [24] report a further bias towards visually salient image features. The authors note that XAI methods predominantly attribute importance to certain image features regardless of whether these features provide target-related information, represent nuisances unrelated to the target, or were entirely absent during model training. Many XAI methods also come in multiple variants, and the criteria for choosing methods and their hyperparameters are often not well justified or documented.
The fundamental limitation of the field, though, is the lack of formal specifications of XAI problems. To ensure the fitness of XAI methods for their intended purposes by design, we argue that the current paradigm of algorithm-first development should be superseded by a requirement-driven XAI development and validation process [see also 29]. Such a process might consist of the following six steps :
@@
It is unreasonable to call a mapping from input features to real numbers an explanation without endowing these numbers with a well-defined formal interpretation [e.g. 62]. As indicated above, different stakeholders, such as ML developers, users (e.g., physicians or patients), and regulators, may intend to use XAI for different purposes associated with different information needs. These needs may concern properties of a given ML model, its training data, a given test input, or combinations of these, and may differ between use cases.
@@
To enable the targeted development of XAI methods tailored to specific purposes, informal information needs communicated by stakeholders need to be translated into formal specifications and requirements, which will inevitably lead to distinct XAI problems. In that sense, “explanation” is understood here as an umbrella term describing the provision of information to stakeholders. We acknowledge that there can be multiple distinct notions of explanation, and thus explanation correctness, depending on the information requested by stakeholders and provided by XAI. We note that, as the correctness of a formalization cannot itself be readily formally validated, it is critical to employ alternatives measures to ensure that it matches stakeholder intentions. To this end, the assessment of stakeholder needs and their subsequent formalization should go hand in hand and is best carried out by seeking consensus in inter-disciplinary and inter-professional teams [c.f., 99].
An example for XAI problem formalization is the work of Karimi et al. [51] on minimal algorithmic recourse, defined as the identification of minimal interventions in model inputs that lead to desired changes in model output. The authors formulate this problem as a constrained optimization over interventions in a given structural causal model (SCM). It is worth noting, though, that counterfactual and recourse explanations may suggest interventions on causal effects of the target or completely non-informative suppressor variables, changes in both of which would have no effect on the target in the real world [94]. More generally, previous work[92, 93, 19] proposes the SAP as a necessary, though not sufficient, requirement for important features in the context of the explanation purposes discussed here.
@@
Given formal problem or requirement specifications, it becomes possible to theoretically analyze existing XAI algorithms with respect to established formal criteria. This has led to the identification of various undesired properties and systematic failure modes of existing XAI methods [e.g., 78, 14]. Kindermans et al. [54], Wilming et al. [94]m, Frye et al. [fryeAsymmetricShapleyValues2020], and Martin and Haufe [57] analyzed popular feature attribution methods and found that many do not fulfill the SAP property in general. Such analyses can help to identify theoretical shortcomings and guide the development of novel, improved methods.
Similarly, it may be possible to devise algorithms that meet formal criteria. The Pattern approach [39] relates a fitted linear model univariately to each individual input feature, thereby avoiding possible misinterpretations due to correlated features. Pattern can consequently be shown to correctly reject suppressor variables in the studied Examples A and B. Recent generalizations such as PatternGAM [26] and PatternLocal [36] extend the Pattern concept to non-linear models and have shown performance gains in empirical benchmarks involving non-linear data [25]. These feature attribution methods fulfill the SAP under well-defined conditions, thereby ensuring that features with high attribution represent sensible starting points, if not solutions, for a variety of popular explanation goals.
@@
Methods that are theoretically and/or empirically validated with respect to given XAI problems or goals should also be theoretically analyzed, quantitatively benchmarked, and algorithmically improved with respect to secondary quality indicators such as robustness, fairness, uncertainty calibration, and alignment with human knowledge in real-world settings. To this end, dedicated benchmarking frameworks such as Quantus [41] can be of use. Moreover, it is crucial to present explanations in ways that are aligned with human cognition and social norms [59]. Note that, although these goals do not necessarily contradict each other, it is an open question whether they can be jointly accomplished in general.. Finally, it can be worthwhile to expand the range of applicability of validated XAI methods. Along these lines, recent work has extended the concept of activation patterns [39] to non-linear and local explanation domains [26, 36].
@@
Just as ML in general, the field of XAI is fast-paced with clever novel methodological developments and empirical validation approaches being introduced each year. Recent advancements in confounder detection [48], generative modeling [44], and causal representation learning [75, 4] promise to address some of the limitations presented here. The systematic formalization and scrutinization of the field of XAI is a wider effort that will eventually make it possible to objectively assess the ability of approaches to solve specific XAI problems. This may lead to XAI-based workflows that can indeed be used to systematically perform quality assurance for ML – and that may eventually find their way into ML production processes and industry standards.